# Inclusive Training Separation and Implicit Knowledge Interaction for Balanced Online Class-Incremental Learning

### v1

@@ 5.3 Ablation Studies
Table 2: Average accuracy with or without using the NCM classifier on Split Mini-ImageNet with various buffer sizes.
@@
BOIL (ours) $\bm{23.9\pm 0.6}$ $\bm{27.0\pm 0.6}$ $\bm{28.5\pm 0.5}$
Impact of Each Component.  Table 3 presents the results of our ablation study on Split Mini-ImageNet with 1k buffer samples. To assess the impact of different components in the BOIL method, we examine performance changes after the removal of individual components: Dual Classifiers (DC), redesigned Proxy-Anchor Loss (PAL), and Momentum Updating (MU), while commonly using the NCM classifier for inference. We first verify the impact of DC by removing DC and MU together, as MU is a backward transfer strategy that depends on DC and cannot function independently. The results, indicated by the reduced AA and increased AI in the ‘w/o DC & MU’ setting, demonstrate the critical role of DC in enhancing plasticity as well as accuracy. This is partly because removing DC also includes discarding its underlying structure using cosine normalization. The comparison between BOIL and ‘w/o MU’ shows that MU further improves both plasticity and accuracy when combined with DC. Removing PAL also results in a decline in both stability and plasticity corresponding to increased AF and AI values in ‘w/o PAL’, leading to a decrease in AA. The ‘w/o all’ setting, which represents vanilla ER with NCM, exhibits the worst overall performance. This indicates that every component is essentially required to obtain the final performance and to maintain a good balance between AF and AI.
@@
Impact of NCM.  From the result of the ‘w/o all’ setting in Table 3, the NCM classifier itself seems to be effective at enhancing stability even when used with vanilla ER. Motivated by this, in Table 2, we further examine whether NCM can similarly enhance performance when combined with competitive state-of-the-art methods on Split Mini-ImageNet with varying buffer sizes. Note that the NCM classifier is already embedded in the SCR method as well as in our BOIL method. As presented in Table 2, our BOIL method still outperforms all the compared methods regardless of whether they use NCM or not. While NCM generally improves classification accuracy, its impact varies depending on the underlying training methods. Specifically, vanilla ER takes the greatest benefit of using NCM, probably because the NCM classifier can mitigate class bias toward new tasks, which is a significant issue in the linear classifier of vanilla ER. In contrast, the performance improvement is less pronounced or occasionally negative in more advanced methods like ER-ACE and PCR. Notably, compared to SCR, which also utilizes NCM in its methodology, BOIL substantially improves the performance with the help of our proposed techniques.
Table 3: Ablation study on Split Mini-ImageNet ($\mathcal{M}$=1k). “w/o all" represents vanilla ER with the NCM classifier.
  Method AA $\uparrow$ AF $\downarrow$ AI $\downarrow$
  BOIL (ours) $\bm{23.9\pm 0.6}$ $9.9\pm 0.7$ $\bm{6.3\pm 8.6}$
w/o DC & MU $18.1\pm 1.3$ $\bm{7.4\pm 0.4}$ $14.8\pm 8.4$
w/o MU $23.2\pm 0.6$ $9.9\pm 0.8$ $6.9\pm 8.9$
w/o PAL $22.8\pm 0.5$ $10.4\pm 0.7$ $7.0\pm 8.4$
w/o all $18.9\pm 1.1$ $8.1\pm 0.9$ $13.1\pm 8.4$
(a) Without DC ($\alpha=1$)
(b) With DC ($\alpha=0$)
Figure 5: Confusion matrix with or without using dual classifiers in Split CIFAR-10 ($\mathcal{M}$ = 1k, $\beta$ = 0, and $m$ = 1). The X-axis is the predicted label and the Y-axis is the ground truth.
Table 4: Balanced performance with or without using dual classifiers in Split CIFAR-10 ($\mathcal{M}$ = 1k, $\beta$ = 0, and $m$ = 1).

### v2

@@ 5.3. Ablation Studies
Table 3. Average accuracy with or without using the NCM classifier on Split Mini-ImageNet with various buffer sizes.
@@
BISON (ours) $\bm{24.2\pm 0.4}$ $\bm{26.9\pm 0.6}$ $\bm{28.7\pm 0.3}$
@@
Impact of NCM. From the result of the ‘w/o all’ setting in Table 2, the NCM classifier itself seems to be effective at enhancing stability even when used with vanilla ER. Motivated by this, in Table 3, we further examine whether NCM can similarly enhance performance when combined with competitive state-of-the-art methods on Split Mini-ImageNet with varying buffer sizes. Note that the NCM classifier is already embedded in the SCR method as well as in our BISON method. As presented in Table 3, our BISON method still outperforms all the compared methods regardless of whether they use NCM or not. While NCM generally improves classification accuracy, its impact varies depending on the underlying training methods. Specifically, vanilla ER takes the greatest benefit of using NCM, probably because the NCM classifier can mitigate class bias toward new tasks, which is a significant issue in the linear classifier of vanilla ER. In contrast, the performance improvement is less pronounced or occasionally negative in more advanced methods like ER-ACE and PCR. Notably, compared to SCR, which also utilizes NCM in its methodology, BISON substantially improves the performance with the help of our proposed techniques.
Figure 5. Performance under different directions of knowledge interaction.
Figure 6. Comparison of execution time on Split CIFAR-10 ($\mathcal{M}$ = 1k).
Directions of Knowledge Interaction. Figure 6 explores alternative directions of knowledge interaction between two classifiers, using Split CIFAR-100 ($\mathcal{M}$ = 1k) and Split CIFAR-10 ($\mathcal{M}$ = 0.5k). Recall that our proposed scheme, denoted as (bs, sb), employs PAL loss for forward transfer (i.e., buffer-to-stream, abbreviated as bs), and applies prototype-level PAF for backward transfer (i.e., stream-to-buffer, abbreviated as sb). We additionally test three possible variants (bs, bs), (sb, sb), and (sb, bs), and compare them with our BISON method (bs, sb). As shown in Figure 6, our proposed direction clearly yields the best AA by achieving an optimal balance between AF and AI. Interestingly, all other alternative directions tend to enhance plasticity but fail to preserve knowledge in the buffer, resulting in lower AAs than ours.
Efficiency Analysis. To study the practical feasibility of our proposed BISON method, we compare the execution times of all baselines on Split CIFAR-10 with a buffer size of 1k. The execution time encompasses both training and inference procedures. As shown in Figure 6, the efficiency of BISON is comparable to ER, SCR, ER-ACE, PCR and LODE, while surpassing them in accuracy. In contrast, methods like GSS, MIR, ASER and ER-DVC, introduce their proposed strategies for buffer management or buffer sample retrieval, which incur significant computational overhead. Similarly, Gdumb involves relatively extensive training with buffer samples over several epochs, which is rare in typical online learning methods, and OCM utilizes rotation augmentation, which also adds to the processing time. Moreover, as presented in Table A4 of Appendix, BISON increases model size by only 0.1% (CIFAR-10) and 1.4% (CIFAR-100), with GPU usage increasing less than 1%. Overall, our proposed method is efficient yet delivers the best prediction performance.
Figure 7. Confusion matrix on Split CIFAR-10 ($\mathcal{M}$ = 0.5k). The X-axis is prediction and the Y-axis is the label. Each off-diagonal white outline represents the top-1 misprediction for each class.
5.4. Quantification of Cross-Task Confusion
Although we have confirmed the effectiveness of each component in our ablation study shown in Table 2, we further conduct an in-depth similar-pair analysis on the benefit of our design.
Metrics. For any class $c$, we define the similar-neighbor set $N(c)$ according to its semantic-similar pair. The global confusion matrix $M\in\mathbb{N}^{C\times C}$ is obtained after training all incremental tasks in a prefixed order and we normalize it along the row to get conditional prediction probabilities: $M_{\mathrm{row}}[c,d]=\mathrm{Pr}(\hat{y}=d|y=c)=\frac{M[c,d]}{\sum_{j}{M[c,j]}+\epsilon}$. Here, we formulate Similar-Confusion at top-1 ($\mathrm{SC@1}$) for per class as $\mathrm{SC@1}(c)=\sum_{d\in N(c)}M_{\mathrm{row}}{[c,d]}$, which reflects proportion of errors that specifically go to the similar neighbors. Moreover, we consider the precision for each class inside the similar pair ($P_{\mathrm{sim}}$), defined as $P_{\mathrm{sim}}(c)=\frac{M_{\mathrm{row}}[c,c]}{M_{\mathrm{row}}[c,c]+\sum_{d\in N(c)}M_{\mathrm{row}}[c,d]}$, measuring discriminability against its most confusable classes.
Protocol. Focusing on the semantic proximity of class labels (WordNet/Miller95) and human intuition (HumanIntuition/PetersonBGR19), we define five semantic-similar pairs for the CIFAR-10 dataset, which comprise {(Cat-Dog), (Deer-Horse), (Automobile-Truck), (Airplane-Ship), (Bird-Frog)}. Then, to construct the incremental tasks, we distribute the classes from these similarity pairs across different tasks, introduced in the following order: {Cat, Deer}, {Dog, Automobile}, {Horse, Airplane}, {Truck, Bird}, and {Ship, Frog}. The results are obtained by making predictions against 10k test samples, 1k pre class after training the last task {Ship, Frog} on Split CIFAR-10 with 0.5k memory buffer samples and all the results are averaged over 3 runs.
Reducing Bias Towards New Tasks. Our first observation is that predictions for the newest classes, Ship and Frog, are reduced when applying our inclusive separation and knowledge interaction. To mitigate the effect of NCM itself on reducing task-recency bias, we apply NCM classifier on both ER and SS-IL during the inference phase. As shown in Figure 7, for ER and SS-IL, most of the off-diagonal white outlines are located in the prediction columns of Frog and Ship, especially the former, where the predictions to Frog are even more than the correct predictions to the label itself (Cat and Deer in ER and Cat in SS-IL). Moreover, our proposed design enhances the precision of predictions for these new classes: the precision for Ship improves from ER’s $\frac{779}{1,534}=50.8\%$ to $\frac{692}{1,122}=61.7\%$, and for Frog from ER’s $\frac{811}{2,221}=36.5\%$ to $\frac{714}{1,324}=53.9\%$, whereas SS-IL achieves $\frac{562}{1,255}=44.8\%$ and $\frac{599}{1,717}=34.9\%$, respectively.
Figure 8. Comparison of similar-confusion at top-1 ($\mathrm{SC}@1$) and precision for each class inside the similar-pair ($P_{sim}$) across tasks on Split CIFAR-10 ($\mathcal{M}$ = 0.5k).
Capturing Semantic Relationship. Figure 7 also indicates that our inclusive separation with knowledge interaction bridges semantically similar classes and preserves cross-task semantic relations, as expected in Figure 3. We follow the main diagonal and focus on each block containing four cells formed by similar-pair. Our BISON method shows more white outlines inside the predefined similar-pair blocks (e.g., CatDog, DeerHorse and Automobile-Truck) than the baseline ER and SS-IL, indicating that residual errors remain within semantic neighbors rather than drifting to other classes. However, this increased correlation is what we intend to achieve for better feature embeddings, as it helps reduce incorrect predictions for unrelated classes, thereby enhancing overall accuracy. As presented in Figure 8, the mass of those confusions decreases ($\mathrm{SC}@1\downarrow$) and the within-neighbor precision increases ($P_{\mathrm{sim}}\uparrow$ ) over tasks, consistently better than ER and SS-IL, i.e., the our BISON method incrementally preserves semantic structure and improves discrimination within the semantic similar-pair.
6. Conclusion
We proposed a novel method for online class-incremental learning, named BISON, to overcome the limitations of existing replay-based methods, which often favor either plasticity or stability. BISON suggests employing dual classifiers to inclusively separate the learning of new samples from the replay of buffer samples, thereby mitigating bias toward new tasks yet effectively consolidating knowledge across different tasks. By introducing implicit knowledge exchange between the dual classifiers, BISON facilitates both forward and backward transfer over the incremental learning process. Empirical results show that BISON achieves the best performance by providing the most balanced learner for OCIL.
Acknowledgements.
This work was supported in part by Institute of Information & communications Technology Planning & Evaluation (IITP) grants funded by the Korea government (MSIT) (No.2022-0-00448, Deep Total Recall: Continual Learning for Human-Like Recall of Artificial Neural Networks), and in part by INHA UNIVERSITY Research Grant.
Supplementary Material: Balanced Online Class-Incremental Learning via Dual Classifiers
In this appendix, we first present a detailed pseudocode of our BISON method, which includes both training and inference processes. Then, we cover more comprehensive experimental results and full implementation details.
Algorithm of Balanced Inclusive Separation for Online iNcremental learning (BISON)
Algorithm 1 shows detailed steps on how BISON trains a neural network $\Theta$ with a given sequence of $T$ tasks, and how it makes a prediction for each test sample after incremental training. For each stream batch $\mathcal{B}^{s}_{\mathcal{D}_{t}}$ at the step $s$, we first randomly retrieve as many buffer samples as there are stream samples in $\mathcal{B}^{s}_{\mathcal{D}_{t}}$ (e.g., 10 samples) (Line 5), then apply data augmentation to both stream and buffer samples, thereby obtaining samples concatenated with their augmented counterparts (Lines 7–9). These original and augmented samples are then fed into the shared feature extractor $h(\cdot~;\Phi)$ to get their corresponding feature vectors (Lines 8–9), and each feature vector is accordingly assigned to either the stream classifier $f(\cdot~;\mathbf{W}_{str})$ or the buffer classifier $f(\cdot~;\mathbf{W}_{buf})$ for the computation of cross-entropy loss terms in Eq. (2) (Lines 10–14). Additionally, the proxy-anchor loss is computed with respect to the buffer features $\mathbf{z}_{\mathcal{M}}$ and class proxies $\mathbf{W}_{str}$ in the stream classifier (Line 15). Based on previously stored label $y^{s-1}_{\mathcal{M}}$, the loss $\mathcal{L}_{Align}$ is calculated between corresponding rows in weights $\mathbf{W}_{str}$ and $\mathbf{W}_{buf}$ (Line 16). The final loss $\mathcal{L}_{BISON}$ is computed and used to update the network parameters (Line 17). To incorporate each stream batch into the buffer, we employ a standard buffer management scheme using reservoir sampling (Line 18). After each training step with a mini-batch, we store the label of buffer batch (Line 21).
After training is completed for each $t$-th task, we perform a testing procedure with test samples from all observed classes in $C_{1:t}$. During inference, we utilize the NCM classifier, where class-wise centroids are computed based on all the samples currently in the buffer (Lines 23–27).
Algorithm 1 Balanced Inclusive Separation for Online iNcremental learning
0: Dataset $\mathcal{D}$; Learning Rate $\lambda$; Adaptive separation smoother $\alpha$; PAL Coefficient $\beta$; PAF Coeffcient $\lambda$.
0: Memory Buffer $\mathcal{M}\xleftarrow{}\{\}$; Network Parameters $\Theta=\{\Phi,\mathbf{W}_{str},\mathbf{W}_{buf},\eta_{str},\eta_{buf},\alpha\}$; Data Augmentation $\textsc{Aug}(\cdot)$.
1:for $t=1$ to $T$ do
2:  /* Training Phase: */
3:  Initialize $\alpha\xleftarrow{}0$
4:  for mini-batch $\mathcal{B}^{s}_{\mathcal{D}_{t}}\sim\mathcal{D}_{t}$ do
5:   $\mathcal{B}^{s}_{\mathcal{M}}\leftarrow\textsc{RandomRetrieval}(\mathcal{M})$ s.t. $|\mathcal{B}^{s}_{\mathcal{M}}|=|\mathcal{B}^{s}_{\mathcal{D}_{t}}|$
6:   for $(\mathbf{x}^{s}_{\mathcal{D}_{t}},y^{s}_{\mathcal{D}_{t}})\in\mathcal{B}^{s}_{\mathcal{D}_{t}},~(\mathbf{x}^{s}_{\mathcal{M}},y^{s}_{\mathcal{M}})\in\mathcal{B}^{s}_{\mathcal{M}}$ do
7:    $\mathbf{x}_{\mathcal{M}}^{aug},\mathbf{x}_{\mathcal{D}_{t}}^{aug}\xleftarrow{}\textsc{Aug}(\mathbf{x}^{s}_{\mathcal{M}}),\textsc{Aug}(\mathbf{x}^{s}_{\mathcal{D}_{t}})$
8:    $\mathbf{z}_{\mathcal{D}}\xleftarrow{}h(\textsc{Concat}([\mathbf{x}^{s}_{\mathcal{D}_{t}},\mathbf{x}_{\mathcal{D}_{t}}^{aug}]);\Phi)$
9:    $\mathbf{z}_{\mathcal{M}}\xleftarrow{}h(\textsc{Concat}([\mathbf{x}^{s}_{\mathcal{M}},\mathbf{x}^{aug}_{\mathcal{M}}]);\Phi)$
10:    $\mathcal{L}_{CE_{str}^{\mathcal{D}}}\xleftarrow{}\textsc{CrossEntropy}(\textsc{Concat}([y^{s}_{\mathcal{D}_{t}},y^{s}_{\mathcal{D}_{t}}]),f(\mathbf{z}_{\mathcal{D}};\mathbf{W}_{str}))$
11:    $\mathcal{L}_{CE_{str}^{\mathcal{M}}}\xleftarrow{}\textsc{CrossEntropy}(\textsc{Concat}([y^{s}_{\mathcal{M}},y^{s}_{\mathcal{M}}]),f(\mathbf{z}_{\mathcal{M}};\mathbf{W}_{str}))$
12:    $\mathcal{L}_{CE_{buf}^{\mathcal{M}}}\xleftarrow{}\textsc{CrossEntropy}(\textsc{Concat}([y^{s}_{\mathcal{M}},y^{s}_{\mathcal{M}}]),f(\mathbf{z}_{\mathcal{M}};\mathbf{W}_{buf}))$
13:    $\alpha\xleftarrow{}\textsc{Sigmoid}(\alpha)$
14:    $\mathcal{L}_{DC}\xleftarrow{}\mathcal{L}_{CE_{str}^{\mathcal{D}}}+\alpha\mathcal{L}_{CE_{str}^{\mathcal{M}}}+(1-\alpha)\mathcal{L}_{CE_{buf}^{\mathcal{M}}}$ // Eq. (2)
15:    $\mathcal{L}_{PAL}\xleftarrow{}\textsc{PAL}(\mathbf{z}_{\mathcal{M}},\textsc{Concat}([y^{s}_{\mathcal{M}},y^{s}_{\mathcal{M}}]),\mathbf{W}_{str})$ // Eq. (3)
16:    $\mathcal{L}_{Align}\xleftarrow{}\textsc{PAF}(\mathbf{W}_{str},\mathbf{W}_{buf},\textsc{Concat}([y^{s-1}_{\mathcal{M}},y^{s-1}_{\mathcal{M}}]))$ // Eq. (4)
17:    $\Theta\xleftarrow{}\Theta+\lambda\nabla_{\Theta}\mathcal{L}_{BISON}$ // Eq. (5)
18:    $\mathcal{M}\xleftarrow{}\textsc{ReservoirUpdate}(\mathcal{M},(\mathbf{x}^{s}_{\mathcal{D}_{t}},y^{s}_{\mathcal{D}_{t}}))$
19:   end for
20:  end for
21:  Store the label $y^{s}_{\mathcal{M}}$
22:  /* Inference Phase: */
23:  for $l\in C_{1:t}$ do
24:   $n_{l}\xleftarrow{}$ number of buffer samples of class $l$
25:   $\bm{\mu}_{l}=\frac{1}{n_{l}}\sum\limits_{(\mathbf{x},y)\in\mathcal{M}}h\left(\mathbf{x};\Phi\right)\cdot\mathbb{I}\left\{{y}=l\right\}$
26:  end for
27:  Given a test sample $\mathbf{x}$, $\hat{y}=\underset{l\in C_{1:t}}{\arg\min}\|h_{\Phi}(\mathbf{x})-\bm{\mu}_{l}\|$ // make a prediction using the NCM classifier
28:end for
Additional Experimental Results
Table A1. Average forgetting at the end of training on three datasets. ‘T’ and ‘F’ indicate two categories: training distinction and feature enhancement, respectively.
Method Split CIFAR-100 Split CIFAR-10 Split Mini-ImageNet
FINE-TUNE 41.4 ±1.3 68.6 ±1.7 35.4 ±0.9
ER 35.8 ±1.0 33.2 ±1.0 33.1 ±1.3 42.9 ±5.0 40.0 ±5.3 34.2 ±7.2 32.3 ±1.4 30.7 ±0.9 31.4 ±1.2
GSS (T) 34.1 ±0.9 32.5 ±1.7 32.7 ±1.0 62.2 ±1.7 56.4 ±3.7 45.2 ±5.7 33.2 ±1.0 32.8 ±1.3 31.4 ±1.3
MIR (T) 33.9 ±1.2 29.9 ±1.5 29.4 ±1.1 40.0 ±3.9 31.5 ±3.4 23.6 ±4.3 29.2 ±1.4 26.6 ±1.6 24.6 ±1.7
A-GEM (T) 43.0 ±1.5 42.1 ±1.1 42.8 ±1.5 67.6 ±3.4 65.5 ±4.2 67.6 ±3.8 37.2 ±0.7 38.3 ±1.0 37.6 ±0.8
Gdumb (T) 15.4 ±0.6 17.2 ±1.1 18.0 ±1.0 23.4 ±4.0 21.6 ±2.6 19.9 ±4.1 13.4 ±0.6 15.9 ±0.7 18.0 ±1.1
SCR (F) 9.8 ±0.8 9.8 ±0.9 9.4 ±0.7 36.8 ±2.0 24.5 ±2.1 20.6 ±2.1 8.1 ±0.6 7.9 ±0.7 7.3 ±0.6
ASER (T) 37.1 ±1.6 33.4 ±1.4 27.9 ±1.4 57.5 ±2.9 50.6 ±4.3 41.4 ±4.1 34.5 ±1.9 31.2 ±1.2 26.4 ±1.4
SS-IL (T) 10.8 ±1.3 11.5 ±0.9 10.7 ±1.4 12.8 ±3.7 10.8 ±2.0 11.7 ±2.0 10.9 ±0.9 9.7 ±1.5 10.3 ±1.2
ER-DVC (T) 30.8 ±1.2 28.0 ±0.8 26.0 ±1.3 32.4 ±4.6 28.8 ±4.6 19.6 ±4.6 27.9 ±1.9 26.9 ±0.9 25.6 ±1.2
ER-ACE (T) 11.5 ±1.0 8.9 ±0.8 8.2 ±0.8 22.4 ±2.3 16.6 ±1.9 12.8 ±3.0 11.0 ±1.6 8.4 ±1.5 7.0 ±1.6
OCM (T, F) 3.7 ±0.4 3.0 ±0.4 2.4 ±0.6 8.2 ±1.6 7.6 ±2.2 5.0 ±1.2 3.3 ±0.5 2.5 ±0.4 2.3 ±0.6
PCR (F) 15.0 ±0.9 11.3 ±1.7 8.3 ±1.2 17.4 ±5.3 12.9 ±2.3 10.1 ±2.2 14.5 ±1.4 11.3 ±0.9 8.9 ±1.0
LODE (T) 15.4 ±1.4 13.3 ±1.0 11.9 ±1.6 14.8 ±1.7 13.2 ±3.4 11.1 ±2.7 14.8 ±1.1 11.8 ±1.4 9.8 ±0.6
BISON (ours) 14.9 ±1.1 9.4 ±0.8 6.9 ±0.5 18.8 ±1.9 15.4 ±1.0 11.5 ±0.9 10.0 ±0.6 7.1 ±0.5 5.6 ±0.5
Table A2. Average intransigence at the end of training on three datasets. ‘T’ and ‘F’ indicate two categories: training distinction and feature enhancement, respectively.
Method Split CIFAR-100 Split CIFAR-10 Split Mini-ImageNet
FINE-TUNE 0 0 0
ER -5.8 ±3.3 -6.4 ±4.4 -7.0 ±3.2 4.9 ±5.1 3.2 ±1.6 4.9 ±5.0 -6.5 ±2.5 -6.8 ±3.0 -5.7 ±3.0
GSS (T) -4.2 ±5.0 -4.7 ±4.9 -4.4 ±5.1 -1.3 ±3.0 -0.6 ±1.7 2.1 ±3.2 -6.2 ±3.9 -7.4 ±3.9 -6.0 ±4.7
MIR (T) -5.3 ±6.2 -3.6 ±6.8 -3.0 ±5.0 8.9 ±6.2 10.8 ±3.4 18.3 ±7.5 -4.2 ±7.6 -2.4 ±6.2 -0.9 ±7.0
A-GEM (T) -1.8 ±2.2 -0.6 ±2.0 -1.9 ±2.6 0.9 ±1.6 3.4 ±3.8 0.9 ±2.8 -1.7 ±2.3 -2.5 ±2.6 -2.5 ±2.6
Gdumb (T) 21.0 ±19.5 13.2 ±20.4 -0.4 ±19.1 36.2 ±22.4 29.1 ±21.2 22.8 ±17.2 18.1 ±17.2 8.9 ±19.2 -4.8 ±19.5
SCR (F) 23.2 ±11.9 21.9 ±10.8 21.3 ±11.3 3.1 ±7.2 6.7 ±5.4 7.7 ±8.0 19.0 ±9.2 17.6 ±9.3 16.9 ±8.7
ASER (T) -9.8 ±4.3 -8.8 ±3.0 -6.9 ±4.3 -1.9 ±2.0 -0.6 ±3.8 0.1 ±4.8 -9.0 ±3.6 -7.8 ±3.5 -6.4 ±4.2
SS-IL (T) 15.9 ±11.6 12.8 ±9.5 14.3 ±8.4 44.4 ±29.7 45.9 ±23.4 42.2 ±29.5 10.6 ±8.3 9.8 ±7.2 9.8 ±7.2
ER-DVC (T) -3.7 ±4.1 -3.6 ±3.5 -3.4 ±3.2 8.2 ±7.7 12.0 ±5.2 15.2 ±6.3 -4.9 ±4.2 -4.5 ±2.8 -4.4 ±3.5
ER-ACE (T) 12.4 ±10.5 12.7 ±11 11.7 ±11.2 16.2 ±15.3 17.1 ±11.7 17.2 ±12.3 9.3 ±8.6 9.4 ±8.3 9.2 ±8.5
OCM (T, F) 38.7 ±16.1 38.4 ±17.6 38.9 ±16.6 39.7 ±32.0 35.2 ±33.4 33.0 ±30.3 30.0 ±15.0 30.6 ±14.8 30.2 ±14.7
PCR (F) 7.3 ±4.4 8.6 ±5.0 9.5 ±6.6 21.9 ±4.4 25.0 ±7.2 25.3 ±7.3 1.9 ±5.1 2.5 ±6.2 4.4 ±7.4
LODE (T) 6.7 ±9.5 6.7 ±8.3 5.7 ±8.9 21.7 ±16.4 16.4 ±12.8 18.7 ±14.3 3.1 ±8.3 3.2 ±7.2 2.6 ±8.1
BISON (ours) 5.3 ±9.7 7.2 ±9.5 6.8 ±11.4 15.0 ±13.8 12.8 ±10.2 14.2 ±14.9 5.9 ±8.7 6.2 ±10.0 5.8 ±9.5
Average Forgetting and Average Intransigence. In addition to the average accuracy shown in Table 1, we also report the corresponding average forgetting (AF) in Table A1 and the corresponding average intransigence (AI) in Table A2 at the end of training, where all measurements are obtained based on 10 independent runs of experiments. As parts of these results (one buffer for each dataset) are also plotted in Figure 4, our BISON method shows the best balance between AF and AI, even though it does not always take the best for both AF and AI at the same time.
Incremental Performance. Figure A1 shows how the performance of each method changes over the incremental training steps, measured by average accuracy after learning each task. Our BISON method almost always maintains the highest performance throughout the entire training process, except for the first few tasks. Although Gdumb (Gdumb/PrabhuTD20) starts with the best performance in Split CIFAR-100 and Split Mini-ImageNet, its performance declines sharply after a few subsequent tasks due to both low plasticity and stability, as reported in Figure 4. Similarly, OCM, which leads for the first two tasks in Split CIFAR-10, also suffers from a notable performance drop, primarily due to its limited ability of knowledge acquisition, reflected in its low plasticity in Figure 4. Meanwhile, PCR demonstrates competitive performance overall in Table 1 and Figure 4, but struggles with initial tasks, exhibiting considerably lower performance compared to other methods. In contrast, BISON starts effectively enough and continues to outperform other methods in subsequent tasks.
Hyperparameter Sensitivity. We conduct a hyperparameter sensitivity analysis on Split CIFAR-10 ($\mathcal{M}$ = 0.2k) and Split Mini-ImageNet ($\mathcal{M}$ = 1k), as shown in Figure A2. All results are averaged over 10 independent runs. For the smaller dataset, Split CIFAR-10, we examine two hyperparameters, $\beta$ and $\lambda$, ranging over $[0.05,0.1,0.15,0.2]$ and $[0.5,1,1.5,2,2.5,3,4]$, respectively. A clear performance plateau appears around $\lambda\in[2,3]$ with $\beta\in[0.05,0.1]$. Along this plateau, varying $\lambda$ causes only minor fluctuations in average accuracy, whereas increasing $\beta$ from 0.1 to 0.2–0.3 consistently degrades performance. The redesigned Proxy-Anchor loss $\mathcal{L}{PAL}$ is computed between the buffer features $\mathbf{z}{\mathcal{M}}$ and the proxies in $\mathbf{W}_{str}$, and its gradients propagate into both the shared backbone and the stream classifier. Consequently, the stream classifier is indirectly influenced by exemplars of old classes. As $\beta$ increases, it enhances the discriminative ability of the stream classifier, but an excessively large $\beta$ disturbs the stability and overall balance, ultimately degrading performance. Overall, Split CIFAR-10 is more sensitive to $\beta$ than to $\lambda$. On Split Mini-ImageNet, due to the limited number of samples in each buffer batch and the larger number of proxies in the stream classifier, stronger values of $\beta$ and $\lambda$ are required to facilitate effective knowledge interaction. Split Mini-ImageNet exhibits a plateau when $\lambda\approx 10$$15$ with moderate $\beta\in[0.2,0.3]$, where performance becomes relatively more sensitive to $\lambda$. Varying $\beta$ within 0.2–0.3 changes accuracy only slightly. We also explore the balance on these two datasets and find that the best balance is achieved near the same regions as the best overall performance ($52.5\pm 0.7\%$ at $(\lambda,\beta)=(3,0.1)$ and $24.2\pm 0.4\%$ at $(10,0.2)$). In conclusion, BISON shows low sensitivity to hyperparameters, as evidenced by the wide performance plateaus. For Split CIFAR-100, which contains the same number of classes, we directly adopt the best settings from Split Mini-ImageNet.
Comparison Under Another Widespread Setting for Feature Enhancement Methods. OCM (OCM/GuoLZ22) and OnPro (OnPro/0001YHZS23) adopt several proprietary data augmentation techniques such as local and global rotations, and they rely on stronger backbones and larger buffer batches. Moreover, OnPro only supports learning in fixed label order and SCR (SCR/MaiLKS21) is also sensitive to the size of the buffer batch. To ensure fairness, we adapt our BISON to their settings and validate the performance on Split CIFAR-100 under varying memory buffer sizes. We employ ResNet-18 (ResNet/HeZRS16) (with 64 filters) and set the size of each buffer batch to 64. The training tasks and label order are fixed, and results are averaged over 15 independently runs. The results of OCM are taken directly from their original paper, while OnPro only reports results for $\mathcal{M}$ = 1k and 2k. Therefore, we mark these results with $\dagger$} and reproduce their method with $\mathcal{M}$ = 5k for comparison. As shown in Table A3, our method BISON consistently achieves the best performance across varying memory buffer sizes, outperforming other feature enhancement methods.
(a) Split CIFAR-100 ($\mathcal{M}$ = 1k)
(b) Split CIFAR-10 ($\mathcal{M}$ = 0.5k)
(c) Split Mini-ImageNet ($\mathcal{M}$ = 1k)
Figure A1. Average accuracy over the incremental training steps.
Figure A2. Impact of Hyperparameter $\lambda$ and $\beta$ on Split CIFAR-10 ($\mathcal{M}$ = 0.2k) and Split Mini-ImageNet($\mathcal{M}$ = 1k).
Table A3. Comparison under another widespread setting for feature enhancement methods on Split CIFAR-100.
Methods Split CIFAR-100
$\mathcal{M}$ = 1000 $\mathcal{M}$ = 2000 $\mathcal{M}$ = 5000
SCR $26.3\pm 0.6$ $31.7\pm 0.4$ $35.3\pm 0.4$
OCM $28.1\pm 0.3$ $35.0\pm 0.4$ $42.4\pm 0.5$